# A Graph-Theoretic Framework for Free-Parafermion Solvability

### v1

@@ Abstract
We show that a quantum spin system has an exact free-parafermion solution if its frustration graph is an oriented indifference graph. Further, we show that if the frustration graph of a model can be dipath oriented via switching operations, then the model is integrable in the sense that there is a family of commuting independent set charges. Additionally, we establish an efficient algorithm for deciding whether this is possible. Our characterisation extends that given for free-fermion solvability. Finally, we apply our results to solve three qudit spin models.
@@ I Introduction
Free fermions provide a fundamental model for understanding various phenomena in condensed-matter physics, quantum field theory, and high-energy physics. Their straightforward mathematical properties allow for precise calculations and predictions of physical phenomena, such as electronic behaviour in materials and band theory. Additionally, certain spin systems can be mapped to free fermions, allowing for exact calculations and insights into these systems. The classical two-dimensional Ising model is an archetypal example of mapping spins to free fermions. Onsager’s exact computation of the free energy [1] can be understood as a consequence of expressing the transfer matrix in terms of bilinears of fermionic operators [2], a process known as the Jordan–Wigner transformation [3]. This transformation allows for the exact computation of many properties, including the spectrum of the quantum Ising spin chain [4].
@@
Our main result shows that if the frustration graph of the model is an oriented indifference graph, then it admits an exact free-parafermion solution (Theorem 1). Further, we show that if the frustration graph of a model can be dipath oriented via switching operations, then the model is integrable in the sense that there is a family of commuting independent set charges (Theorem 2). Additionally, we establish an efficient algorithm for deciding whether an oriented graph can be dipath oriented via switching operations (Theorem 3). Finally, we apply our results to solve three qudit spin models.
This paper is structured as follows. In Section II, we introduce the necessary preliminaries. Then, in Section III, we discuss our main results on free parafermion solvability and integrability. In Section IV, we present our algorithm for deciding dipath orientability via switching operations. We apply our results to solve three qudit spin models in Section V. Finally, we conclude in Section VI with some remarks and open problems.
@@ II.2 Qudit Models
We consider $n$-qudit systems with Hamiltonians written in the generalised Pauli basis
@@
where $V$ is subset of $\mathbb{Z}_{d}^{2n}$ and $h_{v}=b_{v}w_{v}$ for $b_{v}\in\mathbb{R}{\setminus}\{0\}$ and $w_{v}$ an element of $\mathcal{W}_{d}(n)$ determined by the label $v$ in the natural way. Note that the Hamiltonian is not necessarily Hermitian. However, it can be made Hermitian by adding its Hermitian conjugate. We shall restrict our attention to Hamiltonians that satisfy $h_{u}h_{v}=\omega_{d}^{k}h_{v}h_{u}$ with $k$ in $\{-1,0,1\}$ for all $u$ and $v$ in $V$. This condition is always true when $d=3$.
@@
The frustration graph of a Hamiltonian is the oriented graph $G=(V,E)$ with vertex set $V$ and edge set $E=\left\{(u,v)\mid h_{u}h_{v}=\omega_{d}h_{v}h_{u}\right\}$.
@@ II.3 Parafermions
III Parafermion Results
Baxter’s model [17, 20] was solved using a mapping to free parafermions in Ref. [23]. A similar approach was used to identify the free-fermion spectrum of a model that is not solvable via a Jordan–Wigner transformation [9]. This was generalised in Refs. [32, 31] to a specific family of qudit models, and named equipartition indifference models in Ref. [10]. The approach of Ref. [9] can also be generalised to Baxter’s model and the more general $\tau_{2}$ model [41]. Similar to the approach used for free-fermion solutions to many-body models defined on spin-$1/2$ particles [7, 10, 11], this section characterises qudit models that can be solved via a mapping to free parafermions based on a graph-theoretic criterion. Our main result may be stated as follows.
Theorem 1.
Let $H$ be a qudit Hamiltonian with an oriented indifference frustration graph. Then $H$ has an explicit free-parafermion solution.
We begin by introducing a family of commuting charges defined by sets of Hamiltonian terms that commute, or equivalently the independent sets of the frustration graph.
@@ III Parafermion Results
IV Graph-Theoretic Results
@@
Let $G$ be an oriented graph. There is an efficient algorithm to decide whether there exists a dipath-oriented graph in the switching orbit of $G$.
@@ V Applications
The frustration graph of the model with $n=2$ is shown in Fig. 2. The frustration graph is an oriented indifference graph for all $n\in\mathbb{N}$. Hence, Theorem 1 and Theorem 2 apply. By Theorem 1, the model admits an explicit free-parafermion solution, and the single-particle energies are identified with the roots of the independence polynomial. The independence polynomial can be computed efficiently via the algorithm of Ref. [39] or by solving a recurrence relation obtained by applying Proposition 2. By Theorem 2, the model is integrable in the sense that there is a family of commuting independent set charges.
@@ VI Conclusion & Outlook
We have developed a graph-theoretic framework for free-parafermion solvability. Our results show that if the frustration graph of a model is an oriented indifference graph, then it admits an exact free-parafermion solution. Further, we have shown that if the frustration graph of a model can be dipath oriented via switching operations, then the model is integrable in the sense that there is a family of commuting independent set charges. Additionally, we have established an efficient algorithm for determining whether this is possible. Finally, we applied our results to solve three qudit spin models.
@@ Acknowledgements
We thank Catherine Greenhill, Paul Fendley, and Jannis Ruh for helpful discussions. RLM was supported by the ARC Centre of Excellence for Quantum Computation and Communication Technology (CQC2T), project number CE170100012. SJE was supported with funding from the Defense Advanced Research Projects Agency under the Quantum Benchmarking (QB) program under award no. HR00112230007, HR001121S0026, and HR001122C0074 contracts. DRW is supported by the Australian Research Council. AC acknowledges support from EPSRC under agreement EP/T001062/1, and from EU H2020-FETFLAG-03-2018 under grant agreement no. 820495 (AQTION). The views, opinions and/or findings expressed are those of the authors and should not be interpreted as representing the official views or policies of the Department of Defense or the U.S. Government.

### v2

@@ Abstract
We present a graph-theoretic characterisation of when a quantum spin model admits an exact solution via a mapping to free parafermions. Our characterisation is based on the concept of a frustration graph, which represents the commutation relations between Weyl operators of a Hamiltonian. We show that a quantum spin system has an exact free-parafermion solution if its frustration graph is an oriented indifference graph. Further, we show that if the frustration graph of a model can be dipath oriented via switching operations, then the model is integrable in the sense that there is a family of commuting independent set charges. Additionally, we establish an efficient algorithm for deciding whether this is possible. Our characterisation extends that given for free-fermion solvability. Finally, we apply our results to solve three qudit spin models.
@@ I Introduction
Free fermions provide a fundamental model for understanding various phenomena in condensed-matter physics, quantum field theory, and high-energy physics. Their straightforward mathematical properties allow for precise calculations and predictions, for instance, in the electronic structure of materials. Additionally, certain spin systems can be mapped to free fermions, allowing for exact calculations and insights into these systems. The classical two-dimensional Ising model is an archetypal example of mapping spins to free fermions. Onsager’s exact computation of the free energy [1] can be understood as a consequence of expressing the transfer matrix in terms of bilinears of fermionic operators [2], a process known as the Jordan–Wigner transformation [3]. This transformation allows for the exact computation of many properties, including the spectrum of the quantum Ising spin chain [4].
@@
We show that if the frustration graph of a model can be dipath oriented via switching operations, then the model is integrable in the sense that there is a family of commuting independent set charges (Theorem 1). Our main result shows that if the frustration graph of the model is an oriented indifference graph, then it admits an exact free-parafermion solution (Theorem 2). Additionally, we establish an efficient algorithm for deciding whether an oriented graph can be dipath oriented via switching operations (Theorem 3). Finally, we apply our results to solve three qudit spin models.
This paper is structured as follows. In Section II, we introduce the necessary preliminaries. Then, in Section III, we discuss our main results on free-parafermion solvability and integrability. In Section IV, we present our algorithm for deciding dipath orientability via switching operations. We apply our results to solve three qudit spin models in Section V. Finally, we conclude in Section VI with some remarks and open problems.
@@ II.2 Qudit Models
The elements of $\mathcal{W}_{d}(n)$ are called Weyl operators. We consider $n$-qudit systems with Hamiltonians written in the generalised Pauli basis
@@
where $V$ is subset of $\mathbb{Z}_{d}^{2n}$ and $h_{v}=b_{v}w_{v}$ for $b_{v}\in\mathbb{R}{\setminus}\{0\}$ and $w_{v}$ an element of $\mathcal{W}_{d}(n)$ determined by the label $v$ in the natural way. Note that the Hamiltonian is not necessarily Hermitian. However, it can be made Hermitian by adding its Hermitian conjugate. We shall restrict our attention to Hamiltonians that satisfy $h_{u}h_{v}=\omega_{d}^{k}h_{v}h_{u}$ with $k$ in $\{-1,0,1\}$ for all $u$ and $v$ in $V$. This restriction is necessary for our definition of the frustration graph and is always true when $d=2$ or $d=3$.
@@
The frustration graph of a Hamiltonian is the oriented graph $G=(V,E)$ with vertex set $V$ and edge set $E=\left\{(u,v)\mid h_{u}h_{v}=\omega_{d}h_{v}h_{u}\right\}$, where the notation $(u,v)$ denotes a directed edge from vertex $u$ to vertex $v$.
@@ II.3 Parafermions
III Free-Parafermion Solvability and Integrability
In this section, we establish our results on free-parafermion solvability and integrability. We begin by introducing a family of commuting charges defined by sets of Hamiltonian terms that commute, or equivalently the independent sets of the frustration graph.
@@ III Free-Parafermion Solvability and Integrability
Baxter’s model [17, 20] was solved using a mapping to free parafermions in Ref. [23]. A similar approach was used to identify the free-fermion spectrum of a model that is not solvable via a Jordan–Wigner transformation [9]. This was generalised in Refs. [32, 31] to a specific family of qudit models, and named equipartition indifference models in Ref. [10]. The approach of Ref. [9] can also be generalised to Baxter’s model and the more general $\tau_{2}$ model [42]. This is similar to the approach used for free-fermion solutions to many-body models defined on spin-$1/2$ particles [7, 10, 11]. We now characterise qudit models that can be solved via a mapping to free parafermions based on a graph-theoretic criterion. Our main result may be stated as follows.
Theorem 2.
Fix $d\in\mathbb{Z}_{\geq 2}$. Let $H$ be a qudit Hamiltonian with an oriented indifference frustration graph. There exist single-particle energies $\{\varepsilon_{k}\}$ and parafermionic modes $\{\psi_{r,k}\}$ such that
$H=\sum_{k=1}^{\alpha(G)}\sum_{r\in\mathbb{Z}_{d}}\omega_{d}^{r}\varepsilon_{k}\prod_{p=1}^{d}\psi_{r-p,k}.$ (16)
Further, the single-particle energies $\{\varepsilon_{k}\}$ satisfy $Z_{G}\left(-\varepsilon_{k}^{-d}\right)=0$ for all $k\in[\alpha(G)]$.
@@
IV Dipath Orientation via Switching Operations
@@
Let $G$ be an oriented graph. There is an polynomial-time algorithm to decide whether there exists a dipath-oriented graph in the switching orbit of $G$.
@@ V Applications
The frustration graph of the model with $n=2$ is shown in Fig. 2. The frustration graph is an oriented indifference graph for all $n\in\mathbb{N}$. Hence, Theorem 1 and Theorem 2 apply. By Theorem 1, the model is integrable in the sense that there is a family of commuting independent set charges. By Theorem 2, the model admits an explicit free-parafermion solution, and the single-particle energies are identified with the roots of the independence polynomial. The independence polynomial can be computed efficiently via the algorithm of Ref. [39] or by solving a recurrence relation obtained by applying Proposition 2.
@@ VI Conclusion & Outlook
We have developed a graph-theoretic framework for free-parafermion solvability and integrability. Our results show that if the frustration graph of a model is an oriented indifference graph, then it admits an exact free-parafermion solution. Further, we have shown that if the frustration graph of a model can be dipath oriented via switching operations, then the model is integrable in the sense that there is a family of commuting independent set charges. Additionally, we have established an efficient algorithm for determining whether this is possible. Finally, we applied our results to solve three qudit spin models.
@@ Acknowledgements
We thank Paul Fendley, Catherine Greenhill, and Jannis Ruh for helpful discussions. RLM was supported by the ARC Centre of Excellence for Quantum Computation and Communication Technology (CQC2T), project number CE170100012. SJE was supported with funding from the Defense Advanced Research Projects Agency under the Quantum Benchmarking (QB) program under award no. HR00112230007, HR001121S0026, and HR001122C0074 contracts. DRW is supported by the Australian Research Council. AC acknowledges support from EPSRC under agreement EP/T001062/1, and from EU H2020-FETFLAG-03-2018 under grant agreement no. 820495 (AQTION). The views, opinions and/or findings expressed are those of the authors and should not be interpreted as representing the official views or policies of the Department of Defense or the U.S. Government.